### v3

# Monochromatic Translated Product and Answering Sahasrabudhe’s Conjecture

Abstract
This article resolves two related problems in Ramsey theory on the integers. We show that for any finite coloring of the set of natural numbers, there exist numbers $a$ and $b$ for which the configuration $\{a,b,ab,a(b+1)\}$ is monochromatic. By redefining the variables $a=x$ and $ab=y,$ our configurations transforms into $\{x,y,x+y,\frac{y}{x}\}.$ This finding has two main consequences: first, it disproves a conjecture proposed by J. Sahasrabudhe; second, it establishes a quotient version of the long-standing Hindman’s conjecture, which asks for a monochromatic set of the form $\{x,y,x+y,xy\}$.
[…]
[classification=text]
1 Introduction
Arithmetic Ramsey theory studies the existence of monochromatic configurations within any finite coloring of the integers or the natural numbers $\mathbb{N}$. Here, a “coloring” refers to a finite disjoint partition of $\mathbb{N}$, and a set is said to be monochromatic if it is entirely contained in one part of this partition. A collection $\mathcal{F}$ of subsets of $\mathbb{N}$ is said to be partition regular [^1] Although this notion is often referred to as weakly partition regular in the literature, we omit the term “weakly” throughout this paper. if, for every finite coloring of $\mathbb{N}$, there exists a monochromatic member $F\in\mathcal{F}$.
In 1916, I. Schur [8] proved that the pattern $\{a,b,a+b\}$ is partition regular. Applying the map $n\mapsto 2^{n}$, one immediately obtains that $\{a,b,a\cdot b\}$ is also partition regular. These two results are commonly referred to as the additive Schur theorem and the multiplicative Schur theorem, respectively.
For more than a century, it remained unknown whether these two phenomena could be unified in a single statement. That is, whether $\{a,b,a+b,a\cdot b\}$ is also partition regular (see [2]). In 1979, R. L. Graham and N. Hindman independently established such a result for two-colorings (see [4, Section 4]). Subsequently, N. Hindman conjectured that the statement should hold for all finite colorings, a problem now known as the Hindman conjecture. The strongest known partial result toward this conjecture is due to J. Moreira [6], who proved that there exist $a,b\in\mathbb{N}$ such that the set $\{a,ab,a+b\}$ is monochromatic; see also [1] for a concise proof.
In 2017, while investigating asymmetric variants of Schur’s theorem, J. Sahasrabudhe ( see [7, Question 31.]) conjectured that there exist finite colorings of the natural numbers that avoid monochromatic configurations of the form $\{a,b,a(b+1)\}$. In this paper, we disprove this conjecture by showing that every finite coloring of $\mathbb{N}$ necessarily contains such a configuration. In fact, we establish a stronger result: the larger set $\{a,b,ab,a(b+1)\}$ can always be found monochromatic.
On the other hand, by redefining the variables $a=x,ab=y$ our result shows that the pattern $\{x,y,x+y,\frac{y}{x}\}$ is monochromatic. Our result can therefore be viewed as a quotient version of the Hindman conjecture. The technique of our proof builds upon the argument of Moreira [6], who employed van der Waerden’s theorem, whereas we make iterative use of Hindman’s theorem.
The following statement is our main result, which will be proved in the next section.
Theorem 1.
For any finite coloring of $\mathbb{N}$, there exist $a,b\in\mathbb{N}$ such that the set $\{a,b,ab,a(b+1)\}$ is monochromatic.
2 Proof of Theorem 1
Before we proceed to the proof of our Theorem 1, we need to recall some technical definitions.
Definition 1.
  1. 1.
    For any non-empty set $X,$ denote by $\mathcal{P}_{f}(X)$ the set of all non-empty finite subsets of $X.$
  2. 2.
    For any sequence $\langle x_{n}\rangle_{n}$, denote by
    $FS\left(\langle x_{n}\rangle_{n}\right)=\left\{\sum_{n\in H}x_{n}:H\in\mathcal{P}_{f}(\mathbb{N})\right\}.$
  3. 3.
    A set $A\subseteq\mathbb{N}$ is called an IP set if there exists a sequence $\langle x_{n}\rangle_{n}$ such that $FS\left(\langle x_{n}\rangle_{n}\right)\subseteq A.$
  4. 4.
    For any $R\in\mathbb{N},$ a set $A\subseteq\mathbb{N}$ is called an $IP_{R}$ set if there exists a sequence $\langle x_{n}\rangle_{n=1}^{R}$ such that $FS\left(\langle x_{n}\rangle_{n=1}^{R}\right)\subseteq A.$
  5. 5.
    For any $y\in\mathbb{N},$ and $A\subseteq\mathbb{N},$ denote by
    • $y+A=\{n:n+y\in A\};$ and
    • $y^{-1}A=\{n:ny\in A\}.$
To prove our result, we need the following three technical results. The following lemma is a reformulation of the Hindman theorem [3].
Lemma 1.
[5, Lemma 5.19.2.] Any finite coloring of an IPIP set contains a monochromatic IP subset.
The proof of the next two lemmas is standard, but we give the proofs here for the sake of completeness.
Lemma 2.
For any IP set $A\subseteq\mathbb{N}$ and any natural number $y$, there exists another IP set $A^{\prime}\subseteq A$ satisfying $y^{-1}A^{\prime}\subseteq\mathbb{N}$.
Proof.
By passing to a subsequence if necessary, we may assume that AA contains $FS(\langle x_{n}\rangle)$ for some sequence $\langle x_{n}\rangle\subset(y\mathbb{N}+j)$, where $j\in\{0,1,\ldots,y-1\}$.
Now let $F$ be any finite set with $|F|=y$. Then
$x_{F}=\sum_{n\in F}x_{n}\in y\mathbb{N}.$
Thus every sum of exactly $y$ distinct terms from the sequence is divisible by $y$.
Selecting infinitely many distinct elements from the sequence $\langle x_{F}\rangle$ and using them as generators, it follows that AA contains another finite sums set
$A^{\prime}=FS(\langle y_{n}\rangle),$
where each $y_{n}$ is of the form
$y_{n}=yz_{n}$
for some integer $z_{n}$. And hence
$FS(\langle z_{n}\rangle)=y^{-1}A^{\prime}\subseteq\mathbb{N}.$
This completes the proof. ∎
Lemma 3.
For any IPIP set $A\subseteq\mathbb{N}$, there exists $y\in A$ such that
$(-y+A)\cap A$
is also an IPIP set.
Proof.
Let $A=FS(\langle x_{n}\rangle_{n})$ be an IPIP set generated by a sequence $\langle x_{n}\rangle_{n}$, i.e.
$A=\left\{\sum_{n\in F}x_{n}:F\in\mathcal{P}_{f}(\mathbb{N})\right\}.$
Take $y=x_{1}\in A$. Then any finite sum of distinct elements from $(x_{n})_{n\geq 2}$ is of the form $\sum_{n\in F}x_{n}$ with $F\subseteq\{2,3,\dots\}$, and hence
$\sum_{n\in F}x_{n}+y=\sum_{n\in F\cup\{1\}}x_{n}\in A.$
Thus, every element of $FS(x_{n})_{n\geq 2}$ belongs to $(-y+A)\cap A$. Since $FS(x_{n})_{n\geq 2}$ is itself an IPIP set, the intersection $(-y+A)\cap A$ contains an IPIP set, and hence is IPIP. ∎
Now we are in the position to prove our Theorem 1.
Proof of Theorem 1:.
Suppose that $\mathbb{N}$ is finitely colored, say
$\mathbb{N}=\bigcup_{i=1}^{r}A_{i}.$
Without loss of generality, assume that $A_{1}$ is an IPIP set, and set $D_{0}=A_{1}$. By Lemma 3, we can choose an element $y_{1}\in A_{1}$ such that $(-y_{1}+A_{1})\cap A_{1}$ is an IPIP set. Then, by Lemma 2, we can choose an IPIP set
$C_{1}\subset y_{1}^{-1}\big((-y_{1}+A_{1})\cap A_{1}\big),$
which is a subset of $\mathbb{N}$. Finally, by Lemma 1, there exists $i_{1}\in\{1,2,\ldots,r\}$ such that
$D_{1}=C_{1}\cap A_{i_{1}}=y_{1}^{-1}\big((-y_{1}+A_{1})\cap A_{1}\big)\cap A_{i_{1}}$
is an IPIP set.
Using Lemmas 2 and 3, choose $y_{2}\in D_{1}$ such that
$C_{2}\subseteq y_{2}^{-1}\big((-y_{2}+D_{1})\cap D_{1}\big)\subset\mathbb{N}$
is an IPIP set. Then, by Lemma 1, there exists $i_{2}\in\{1,2,\ldots,r\}$ such that $D_{2}=C_{2}\cap A_{i_{2}}$ is an IPIP set. Proceeding inductively, for each $n\in\mathbb{N}$, choose $y_{n}\in D_{n-1}$ and $i_{n}\in\{1,2,\ldots,r\}$ such that the set
$D_{n}=y_{n}^{-1}\big((-y_{n}+D_{n-1})\cap D_{n-1}\big)\cap A_{i_{n}}$
is an IPIP set.
By the Pigeonhole Principle, there exists $k\in\{1,2,\ldots,r\}$ such that $A_{i_{j}}=A_{i_{n}}=A_{k}$ for some indices $j<n$. Now choose $x\in D_{n}\subseteq A_{i_{n}}=A_{k}$.
From the construction in (2), we have
$xy_{n}\in D_{n-1}\subseteq y_{n-1}^{-1}D_{n-2}\subseteq\cdots\subseteq y_{n-1}^{-1}\cdots y_{j+1}^{-1}D_{j}.$
Hence,
$xy_{n}\cdots y_{j+1}\in D_{j}\subseteq A_{i_{j}}=A_{k}.$
Moreover, since $y_{n}\in D_{n-1}$, applying the same argument gives
$y_{n}\cdots y_{j+1}\in D_{j}\subseteq A_{i_{j}}=A_{k}.$
Thus, both $y_{n}\cdots y_{j+1}$ and $xy_{n}\cdots y_{j+1}$ lie in the same color class $A_{k}$.
Again, observe that
$xy_{n}+y_{n}\in D_{n-1},$
so by the same argument as above,
$(xy_{n}+y_{n})\,y_{n-1}\cdots y_{j+1}\in D_{j}\subseteq A_{i_{j}}.$
Now define
  1. 1.
    $b=x$, and
  2. 2.
    $a=y_{n}\cdots y_{j+1}$.
Then $a,b\in A_{k}$, and
$xy_{n}\cdots y_{j+1}=ab\in A_{k},\quad\text{and}\quad(xy_{n}+y_{n})\,y_{n-1}\cdots y_{j+1}=a(b+1)\in A_{k}.$
Hence, the set
$\{a,b,ab,a(b+1)\}\subseteq A_{k}$
is monochromatic.
This completes the proof.
Acknowledgments
The author is thankful to the referees for their comments on the previous draft of this paper and also for the proof of the Lemma 2. This paper is supported by NBHM postdoctoral fellowship with reference no: 0204/27/(27)/2023/R & D-II/11927.